Many casinos have a game called the Big Six Money Wheel, which has 54 slots in which are displayed a Joker, the casino logo, and various dollar amounts, as shown in the table at the top of the next column. Players may bet on the Joker, the casino logo, or one or more dollar denominations. The wheel is spun and if the wheel stops on the same place as the player's bet, the player wins that amount for each dollar bet.
Denomination Number of slots
$40 (Joker) 1
$40 (Casino logo) 1
$20 2
$10 4
$5 7
$2 15
$1 24
If a player bets $4 on the Casino logo denomination, find the player's expectation. (Round your answer to two decimal places.)

Answer by Boreal(15235)

There is a 1/54 probability of landing on the Casino logo
E(X)=(1/54)*$40-(53/54)*$4
=40/54-212/54
=-172/54 or minus $3.19